OpenStudy (anonymous):

am i correct? division (3x^(2)-12x)/(2x-7)÷(x-4)/(2x-7) answer = 3x/2-3/4-21/4(2x-7) am i right?

hero (hero):

\[\frac{3x^2 - 12x}{2x - 7} \div \frac{x - 4}{2x - 7}\] You have to reciprocate the second fraction then multiply: \[\frac{3x^2 - 12x}{2x - 7} \times \frac{2x - 7}{x - 4}\] Also, you have to cancel factors of one: \[\frac{3x\cancel{(x - 4)}}{\cancel{2x - 7}} \times \frac{\cancel{2x - 7}}{\cancel{x - 4}}\]

OpenStudy (anonymous):

so the answer is 3x? im not good at math sorry lol

hero (hero):

Yes, but you should learn how to do these.

OpenStudy (anonymous):

yes i know im trying to learn thanks

hero (hero):

khan academy can teach you